OpenStudy (anonymous):

if 3+xi=(x+1)+2i what is the value of x?

OpenStudy (mayankdevnani):

compare the terms first and then simplify for `x`

OpenStudy (mayankdevnani):

like in this one, \[\large \bf \mathbb{R}~part=\mathbb{R}~part\] and \[\large \bf Imaginary~part=Imaginary~part\]

OpenStudy (mayankdevnani):

so, we get \[\large \bf 3=x+1\] (Equating Real part) \[\large \bf x=2\] (Equating Imaginary part)

OpenStudy (mayankdevnani):

hope you understand ! @halghafilee

OpenStudy (loser66):